# Conformal prediction with local weights: randomization enables robust guarantees

### v4

@@ Abstract
In this work, we consider the problem of building distribution-free prediction intervals with finite-sample conditional coverage guarantees. Conformal prediction (CP) is an increasingly popular framework for building prediction intervals with distribution-free guarantees, but these guarantees only ensure marginal coverage: the probability of coverage is averaged over a random draw of both the training and test data, meaning that there might be substantial undercoverage within certain subpopulations. Instead, ideally we would want to have local coverage guarantees that hold for each possible value of the test point’s features. While the impossibility of achieving pointwise local coverage is well established in the literature, many variants of conformal prediction algorithm show favorable local coverage properties empirically. Relaxing the definition of local coverage can allow for a theoretical understanding of this empirical phenomenon.
We aim to bridge this gap between theoretical validation and empirical performance by proving achievable and interpretable guarantees for a relaxed notion of local coverage. Building on the localized CP method of Guan, (2023) and the weighted CP framework of Tibshirani et al., (2019), we propose a new method, randomly-localized conformal prediction (RLCP), which returns prediction intervals that are not only marginally valid but also achieve a relaxed local coverage guarantee and guarantees under covariate shift. Through a series of simulations and real data experiments, we validate these coverage guarantees of RLCP while comparing it with the other local conformal prediction methods.
@@ 1 Introduction
Consequently, we are interested in finding a meaningful compromise between test-conditional coverage (2), which cannot be achieved in a distribution-free regime (aside from settings where $X$ is discrete), and marginal coverage (1), which can be achieved distribution-free but is too weak for many practical applications. In particular, we might ask for coverage to hold when we condition, not on a specific value of $X_{n+1}$, but instead in a weaker sense, e.g., for any $X_{n+1}$ within some neighborhood—for instance, $90\%$ coverage may not be guaranteed for patients who are exactly 25.75 years old, but perhaps could be guaranteed to hold on average over any 5-year age range. To formalize this notion mathematically, let’s consider a collection ${\mathcal{B}}$ of subsets of feature space, $B\subseteq{\mathcal{X}}$, such that we seek valid coverage guarantees if the test feature lies in $B$, for each $B\in{\mathcal{B}}$ (for instance, each $B$ might be a subset of patients defined by a particular 5-year age range). This can be expressed as the following desired guarantee:
@@
Different choices of the collection of subsets ${\mathcal{B}}$ represent different coverage goals for $\widehat{C}_{n}(X_{n+1})$. For example, if we take =${\mathcal{B}}=\{{\mathcal{X}}\}$, then the coverage guarantee (3) is simply the marginal coverage guarantee (1). At the other extreme, if =${\mathcal{B}}=\{\{x\}:x\in{\mathcal{X}}\}$ contains all the singleton sets, then (3) becomes equivalent to the (impossible) pointwise test-conditional coverage guarantee (2) (and more broadly, the hardness result in Barber et al., (2021) states that it is impossible to achieve such a guarantee whenever the VC-dimension of ${\mathcal{B}}$ is too large). Therefore, to capture a notion of coverage that lies in between the two, we are interested in collections ${\mathcal{B}}$ that lie somewhere in between these two extremes—for instance, if ${\mathcal{X}}=\mathbb{R}^{d}$, we might consider all open balls of a certain radius.
@@ 1.1 Background: conformal prediction, local coverage, and covariate shift
We will next give some background on the conformal prediction methodology, and prior work that aims to achieve approximate local guarantees for conformal prediction (some of which are similar to (3) or to (5) in flavor).
Conformal inference (CP) (Vovk et al.,, 2005) is an extremely popular framework for calibrating prediction uncertainty, and is used to build prediction intervals satisfying distribution-free marginal validity as in (1). A common variant of this methodology is the split conformal prediction procedure (Papadopoulos et al.,, 2002), which we briefly define now. Assume we have two datasets available $\{(X_{i}^{\textnormal{pre}},Y_{i}^{\textnormal{pre}})\}_{i=1,\dots,n_{\textnormal{pre}}}$, a dataset used for pretraining, and $\{(X_{i},Y_{i})\}_{i=1,\dots,n}$, the holdout set which will be used for calibration. [^1] In practice, of course, we would simply partition a single available dataset into two disjoint subsets—but it will be convenient here to use the notation of two separate datasets. Suppose that we fit a model on $\{(X_{i}^{\textnormal{pre}},Y_{i}^{\textnormal{pre}})\}_{i=1,\dots,n_{\textnormal{pre}}}$, and obtain a conformity score function $s(x,y)$ that returns a large value if the point $(x,y)$ appears unusual (does not “conform”) relative to the trends observed in the data. For example, if we run a regression procedure to train a model $\widehat{f}(x)$ (which predicts $Y$ given $X=x$), we might choose the commonly-used residual score $s(x,y)=|y-\widehat{f}(x)|$, which returns a large value if the value of $y$ is not near the predicted value $\widehat{f}(x)$. The dataset $\{(X_{i},Y_{i})\}_{i=1,\dots,n}$, which contains $n$ many data points that were not used for training $\widehat{f}$, is then used for calibrating a threshold for the score function: defining $\widehat{q}$ as the $(1-\alpha)(1+1/n)$ quantile of the holdout scores $\{s(X_{i},Y_{i})\}_{i=1,\dots,n}$, we return the prediction interval [^2] We use the terminology “prediction interval” throughout the paper as this is most familiar in statistical settings where often we have a real-valued response (i.e., ${\mathcal{Y}}=\mathbb{R}$), but it might be more correct to say “prediction set”, since even when ${\mathcal{Y}}=\mathbb{R}$ the resulting set may not necessarily consist of a single interval.
@@
This method is guaranteed to satisfy distribution-free marginal coverage as in (1)—but what if we would like to achieve (relaxed) test-conditional coverage and/or coverage with respect to (some classes of) covariate shift? If our goal is to achieve test-conditional coverage relative to a single known subset $B\subseteq{\mathcal{X}}$, we can simply run split conformal prediction using only holdout points $i$ with $X_{i}\in B$ to calibrate $\widehat{q}$. Similarly, if our goal is to achieve coverage under covariate for a single known shifted distribution ${\widetilde{P}}_{X}$, the weighted CP method of Tibshirani et al., (2019) uses a weighted quantile, with weights reflecting the shift from $P_{X}$ to ${\widetilde{P}}_{X}$, to define $\widehat{q}$ in a way that achieves the desired coverage guarantee: the threshold $\widehat{q}$ in (6) is replaced with a weighted quantile, $\widehat{q}=\textnormal{Quantile}_{1-\alpha}\left(\sum_{i=1}^{n}w_{i}\delta_{s(X_{i},Y_{i})}+w_{n+1}\delta_{+\infty}\right)$, where weights $w_{i}$ are chosen to correct for the covariate shift; for example, if $P_{X}$ has density $f_{X}$ and ${\widetilde{P}}_{X}$ has density $\widetilde{f}_{X}$, we choose $w_{i}\propto\widetilde{f_{X}}(X_{i})/f_{X}(X_{i})$.
@@ 1.2 Our contributions
Of course, all of these aims could again be achieved by simply returning the uninformative interval $\widehat{C}_{n}(X_{n+1})\equiv{\mathcal{Y}}$—so our goal is to provide a $\widehat{C}_{n}$ that satisfies aims (A1)(A3) without being too conservative, i.e., producing an interval that is as narrow as possible. We will see that RLCP provides a tool that can allow us to achieve these aims without being overly conservative.
@@
The remainder of the paper is organized as follows. We will begin by reviewing the literature on framework for locally-weighted conformal prediction methods in Section 2, giving details on some existing methods and their shortcomings in order to gain motivation for our proposed method RLCP. Then in Section 3, we formally introduce RLCP, and in Section 4, we present theoretical results towards establishing that RLCP satisfies the desired properties (A1), (A2), (A3). In Section 5, we demonstrate the performance of RLCP (and compare it with these other locally-weighted CP methods described in Section 2) on both simulated and real data. Finally, we conclude with a discussion in Section 6, where we compare our method with existing literature and examine the role that randomization plays in our method. Some proofs and some details of implementing the methods are deferred to the Appendix.
@@ 2.1 An initial approach: baseLCP
For the Gaussian kernel and box kernel described above (as well as other natural choice of $H$), the choice of bandwidth $h$ controls the extent of localization. At a high level, for larger values of $h$ the method behaves similarly to split conformal, while for smaller $h$ the method becomes more localized, and can offer better empirical local coverage. However, if we choose $h$ too small, performance will suffer due to low effective sample size—the quantile $\widehat{q}(x)$ may even be estimated as +$\infty$+ ∞. We will return later on to the question of choosing a bandwidth $h$ that appropriately addresses this tradeoff.
@@ 2.2 Calibrating to recover marginal coverage: calLCP
is the set of possible sums of weights that might occur in the calculation of the discrete CDF inside the quantile.
@@ 2.3 Another look at baseLCP: achieving a different coverage guarantee
Now we return to the baseLCP method. How does constructing the “synthetic sample” ${\widetilde{X}}_{n+1}$ relate to this procedure? It turns out that baseLCP is actually related to running weighted conformal prediction (Tibshirani et al.,, 2019) on the covariate shift problem defined by taking ${\widetilde{P}}_{X}=P_{X}\circ H(\cdot,X_{n+1})$. Specifically, if we define
@@
This interpretation leads us towards the RLCP method. In baseLCP, using the kernel $H(\cdot,X_{n+1})$ centered at the true test point $X_{n+1}$, leads to a threshold $\widehat{q}_{1-\alpha}(X_{n+1})$ that provides coverage for a synthetic sample ${\widetilde{X}}_{n+1}$—that is, for data points that are similar to $X_{n+1}$, but not necessarily for $X_{n+1}$ itself. For RLCP, then, we will instead use a kernel that is not centered at $X_{n+1}$, but will aim to achieve coverage for $X_{n+1}$ itself.
3 Randomly-localized conformal prediction
From our discussion above, we saw that baseLCP constructs the threshold $\widehat{q}_{1-\alpha}(X_{n+1})$ with weights defined by a kernel centered at $X_{n+1}$, which then leads to coverage for a “synthetic sample” ${\widetilde{X}}_{n+1}$ that is sampled near $X_{n+1}$. To define randomly-localized conformal prediction (RLCP), we will reverse this construction: we will define the threshold using weights centered at a random ${\widetilde{X}}_{n+1}$ sampled near $X_{n+1}$, and will see later on that this leads to marginal coverage (and also approximate local coverage) at $X_{n+1}$.
@@ 3 Randomly-localized conformal prediction
The only difference between RLCP and baseLCP, then, is that the weights are given by the $\widetilde{w}_{i}$’s, computed with the kernel $H$ centered at ${\widetilde{X}}_{n+1}$, rather than the original $w_{i}$’s, which were computed with $H$ centered at $X_{n+1}$.
@@ 3.1 The role of ${\widetilde{X}}_{n+1}$
Before proceeding, we pause to discuss ${\widetilde{X}}_{n+1}$ a bit more. First, observe that the randomness of ${\widetilde{X}}_{n+1}$ means that the resulting prediction interval is no longer a deterministic function of the data—it is randomized because the threshold $\widehat{q}_{1-\alpha}(X_{n+1},{\widetilde{X}}_{n+1})$ is computed with the kernel $H$ centered at a random point ${\widetilde{X}}_{n+1}$ (hence the term “randomly-localized” in the name of the method). To compare again with baseLCP,
  • For baseLCP, the weights used for computing the threshold $\widehat{q}_{1-\alpha}(X_{n+1})$ are calculated by centering the kernel $H$ at $X_{n+1}$, and can lead to coverage for a “synthetic sample” ${\widetilde{X}}_{n+1}$, drawn from $P_{X}\circ H(\cdot,X_{n+1})$. In other words, we can think of the original test point $X_{n+1}$ as a prototype: coverage will hold relative to the reweighted feature distribution $P_{X}\circ H(\cdot,X_{n+1})$, which generates new “synthetic samples” ${\widetilde{X}}_{n+1}$ which are potentially very similar to this prototype—but coverage may fail for $X_{n+1}$ itself, i.e., marginal coverage may fail to hold, as observed by Guan, (2023).
  • For RLCP, the weights used for computing the threshold $\widehat{q}_{1-\alpha}(X_{n+1},{\widetilde{X}}_{n+1})$ are calculated by centering the kernel $H$ at a random point ${\widetilde{X}}_{n+1}$ that was drawn near $X_{n+1}$, and (as we will see below) will lead to coverage for $X_{n+1}$ itself. We can now think of ${\widetilde{X}}_{n+1}$ as a “synthetic prototype”, and conditionally, the test sample $X_{n+1}$ is likely to lie near to this prototype ${\widetilde{X}}_{n+1}$. More concretely, we will see that the conditional distribution of $X_{n+1}$ given ${\widetilde{X}}_{n+1}$ is equal to the reweighted feature distribution $P_{X}\circ H(\cdot,{\widetilde{X}}_{n+1})$, which (for a strongly localizing kernel $H$) places most of its mass near the “synthetic prototype” ${\widetilde{X}}_{n+1}$.

### v5

@@ Abstract
In this work, we consider the problem of building distribution-free prediction intervals with finite-sample conditional coverage guarantees. Conformal prediction (CP) is an increasingly popular framework for building such intervals with distribution-free guarantees, but these guarantees only ensure marginal coverage: the probability of coverage is averaged over both the training and test data, meaning that there might be substantial undercoverage within certain subpopulations. Instead, ideally we would want to have local coverage guarantees that hold for each possible value of the test point’s features. While the impossibility of achieving pointwise local coverage is well established in the literature, many variants of conformal prediction algorithm show favourable local coverage properties empirically. Relaxing the definition of local coverage can allow for a theoretical understanding of this empirical phenomenon.
We propose randomly-localized conformal prediction (RLCP), a method that builds on localized CP and weighted CP techniques to return prediction intervals that are not only marginally valid but also offer relaxed local coverage guarantees and validity under covariate shift. Through a series of simulations and real data experiments, we validate these coverage guarantees of RLCP while comparing it with the other local conformal prediction methods.
Keywords: distribution free prediction intervals, conformal prediction, local coverage guarantees, localized conformal prediction, weighted conformal prediction, distribution shift
@@ 1 Introduction
Consequently, we are interested in finding a meaningful compromise between test-conditional coverage (2), which cannot be achieved in a distribution-free regime (aside from settings where $X$ is discrete), and marginal coverage (1), which can be achieved distribution-free but is too weak for many practical applications. In particular, we might ask for coverage to hold when we condition, not on a specific value of $X_{n+1}$, but instead in a weaker sense, e.g., for any $X_{n+1}$ within some neighbourhood—for instance, $90\%$ coverage may not be guaranteed for patients who are exactly 25.75 years old, but perhaps could be guaranteed to hold on average over any 5-year age range. To formalize this notion mathematically, let’s consider a collection ${\mathcal{B}}$ of subsets of feature space, $B\subseteq{\mathcal{X}}$, such that we seek valid coverage guarantees if the test feature lies in $B$, for each $B\in{\mathcal{B}}$ (for instance, each $B$ might be a subset of patients defined by a particular 5-year age range). This can be expressed as the following desired guarantee:
@@
Different choices of the collection of subsets ${\mathcal{B}}$ represent different coverage goals for $\widehat{C}_{n}(X_{n+1})$. For example, if we take =${\mathcal{B}}=\{{\mathcal{X}}\}$, then the coverage guarantee (3) is simply the marginal coverage guarantee (1). At the other extreme, if =${\mathcal{B}}=\{\{x\}:x\in{\mathcal{X}}\}$ contains all the singleton sets, then (3) becomes equivalent to the (impossible) pointwise test-conditional coverage guarantee (2) (and more broadly, the hardness result in Barber et al., (2021) states that it is impossible to achieve such a guarantee whenever the Vapnik–Chervonenkis (VC)-dimension of ${\mathcal{B}}$ is too large). Therefore, to capture a notion of coverage that lies in between the two, we are interested in collections ${\mathcal{B}}$ that lie somewhere in between these two extremes—for instance, if ${\mathcal{X}}=\mathbb{R}^{d}$, we might consider all open balls of a certain radius.
@@ 1.1 Background: conformal prediction, local coverage, and covariate shift
We will next give some background on the conformal prediction methodology, and prior work that aims to achieve approximate local guarantees for conformal prediction (some of which are similar to (3) or to (5) in flavour).
Conformal inference (CP) (Vovk et al.,, 2005) is an extremely popular framework for calibrating prediction uncertainty, and is used to build prediction intervals satisfying distribution-free marginal validity as in (1). A common variant of this methodology is the split conformal prediction procedure (Papadopoulos et al.,, 2002), which we briefly define now. Assume we have two datasets available $\{(X_{i}^{\textnormal{pre}},Y_{i}^{\textnormal{pre}})\}_{i=1,\dots,n_{\textnormal{pre}}}$, a dataset used for pretraining, and $\{(X_{i},Y_{i})\}_{i=1,\dots,n}$, the holdout set which will be used for calibration. [^1] In practice, of course, we would simply partition a single available dataset into two disjoint subsets—but it will be convenient here to use the notation of two separate datasets. Suppose that we fit a model on $\{(X_{i}^{\textnormal{pre}},Y_{i}^{\textnormal{pre}})\}_{i=1,\dots,n_{\textnormal{pre}}}$, and obtain a conformity score function $s(x,y)$ that returns a large value if the point $(x,y)$ appears unusual (does not “conform”) relative to the trends observed in the data. For example, if we run a regression procedure to train a model $\widehat{f}(x)$ (which predicts $Y$ given $X=x$), we might choose the commonly used residual score $s(x,y)=|y-\widehat{f}(x)|$, which returns a large value if the value of $y$ is not near the predicted value $\widehat{f}(x)$. The dataset $\{(X_{i},Y_{i})\}_{i=1,\dots,n}$, which contains $n$ many data points that were not used for training $\widehat{f}$, is then used for calibrating a threshold for the score function: defining $\widehat{q}$ as the $(1-\alpha)(1+1/n)$ quantile of the holdout scores $\{s(X_{i},Y_{i})\}_{i=1,\dots,n}$, we return the prediction interval [^2] We use the terminology “prediction interval” throughout the paper as this is most familiar in statistical settings where often we have a real-valued response (i.e., ${\mathcal{Y}}=\mathbb{R}$), but it might be more correct to say “prediction set”, since even when ${\mathcal{Y}}=\mathbb{R}$ the resulting set may not necessarily consist of a single interval.
@@
This method is guaranteed to satisfy distribution-free marginal coverage as in (1)—but what if we would like to achieve (relaxed) test-conditional coverage and/or coverage with respect to (some classes of) covariate shift? If our goal is to achieve test-conditional coverage relative to a single known subset $B\subseteq{\mathcal{X}}$, we can simply run split conformal prediction using only holdout points $i$ with $X_{i}\in B$ to calibrate $\widehat{q}$. Similarly, if our goal is to achieve coverage under covariate for a single known shifted distribution ${\widetilde{P}}_{X}$, the Weighted Conformal Prediction (WCP) method of Tibshirani et al., (2019) uses a weighted quantile, with weights reflecting the shift from $P_{X}$ to ${\widetilde{P}}_{X}$, to define $\widehat{q}$ in a way that achieves the desired coverage guarantee: the threshold $\widehat{q}$ in (6) is replaced with a weighted quantile, $\widehat{q}=\textnormal{Quantile}_{1-\alpha}\left(\sum_{i=1}^{n}w_{i}\delta_{s(X_{i},Y_{i})}+w_{n+1}\delta_{+\infty}\right)$, where weights $w_{i}$ are chosen to correct for the covariate shift; for example, if $P_{X}$ has density $f_{X}$ and ${\widetilde{P}}_{X}$ has density $\widetilde{f}_{X}$, we choose $w_{i}\propto\widetilde{f_{X}}(X_{i})/f_{X}(X_{i})$.
@@ 1.2 Our contributions
Of course, all of these aims could again be achieved by simply returning the uninformative interval $\widehat{C}_{n}(X_{n+1})\equiv{\mathcal{Y}}$—so our goal is to provide a $\widehat{C}_{n}$ that satisfies aims (A1)(A3) without being too conservative, i.e., producing an interval that is as narrow as possible. We will see that randomly localized conformal prediction (RLCP) provides a tool that can allow us to achieve these aims without being overly conservative.
@@
The remainder of the paper is organized as follows. We will begin by reviewing the literature on framework for locally weighted conformal prediction methods in Section 2, giving details on some existing methods and their shortcomings in order to gain motivation for our proposed method RLCP. Then in Section 3, we formally introduce RLCP, and in Section 4, we present theoretical results towards establishing that RLCP satisfies the desired properties (A1), (A2), (A3). In Section 5, we demonstrate the performance of RLCP (and compare it with these other locally weighted CP methods described in Section 2) on both simulated and real data. Finally, we conclude with a discussion in Section 6, where we compare our method with existing literature and examine the role that randomization plays in our method. Some proofs and some details of implementing the methods are deferred to the Appendix.
@@ 2.1 An initial approach: baseLCP
For the Gaussian kernel and box kernel described above (as well as other natural choice of $H$), the choice of bandwidth $h$ controls the extent of localization. At a high level, for larger values of $h$ the method behaves similarly to split conformal, while for smaller $h$ the method becomes more localized, and can offer better empirical local coverage. However, if we choose $h$ too small, performance will suffer due to low effective sample size—the quantile $\widehat{q}(x)$ may even be estimated as +$\infty$+ ∞. We will return later on to the question of choosing a bandwidth $h$ that appropriately addresses this trade-off.
@@ 2.2 Calibrating to recover marginal coverage: calLCP
is the set of possible sums of weights that might occur in the calculation of the discrete Cumulative Distribution Fucntion (CDF) inside the quantile.
@@ 2.3 Another look at baseLCP: achieving a different coverage guarantee
Now we return to the baseLCP method. How does constructing the “synthetic sample” ${\widetilde{X}}_{n+1}$ relate to this procedure? It turns out that baseLCP is actually related to running WCP (Tibshirani et al.,, 2019) on the covariate shift problem defined by taking ${\widetilde{P}}_{X}=P_{X}\circ H(\cdot,X_{n+1})$. Specifically, if we define
@@
This interpretation leads us towards the RLCP method. In baseLCP, using the kernel $H(\cdot,X_{n+1})$ centred at the true test point $X_{n+1}$, leads to a threshold $\widehat{q}_{1-\alpha}(X_{n+1})$ that provides coverage for a synthetic sample ${\widetilde{X}}_{n+1}$—that is, for data points that are similar to $X_{n+1}$, but not necessarily for $X_{n+1}$ itself. For RLCP, then, we will instead use a kernel that is not centred at $X_{n+1}$, but will aim to achieve coverage for $X_{n+1}$ itself.
3 Randomly localized conformal prediction
From our discussion above, we saw that baseLCP constructs the threshold $\widehat{q}_{1-\alpha}(X_{n+1})$ with weights defined by a kernel centred at $X_{n+1}$, which then leads to coverage for a “synthetic sample” ${\widetilde{X}}_{n+1}$ that is sampled near $X_{n+1}$. To define RLCP, we will reverse this construction: we will define the threshold using weights centred at a random ${\widetilde{X}}_{n+1}$ sampled near $X_{n+1}$, and will see later on that this leads to marginal coverage (and also approximate local coverage) at $X_{n+1}$.
@@ 3 Randomly localized conformal prediction
The only difference between RLCP and baseLCP, then, is that the weights are given by the $\widetilde{w}_{i}$’s, computed with the kernel $H$ centred at ${\widetilde{X}}_{n+1}$, rather than the original $w_{i}$’s, which were computed with $H$ centred at $X_{n+1}$.
@@ 3.1 The role of ${\widetilde{X}}_{n+1}$
Before proceeding, we pause to discuss ${\widetilde{X}}_{n+1}$ a bit more. First, observe that the randomness of ${\widetilde{X}}_{n+1}$ means that the resulting prediction interval is no longer a deterministic function of the data—it is randomized because the threshold $\widehat{q}_{1-\alpha}(X_{n+1},{\widetilde{X}}_{n+1})$ is computed with the kernel $H$ centred at a random point ${\widetilde{X}}_{n+1}$ (hence the term “randomly localized” in the name of the method). To compare again with baseLCP,
  • For baseLCP, the weights used for computing the threshold $\widehat{q}_{1-\alpha}(X_{n+1})$ are calculated by centring the kernel $H$ at $X_{n+1}$, and can lead to coverage for a “synthetic sample” ${\widetilde{X}}_{n+1}$, drawn from $P_{X}\circ H(\cdot,X_{n+1})$. In other words, we can think of the original test point $X_{n+1}$ as a prototype: coverage will hold relative to the reweighted feature distribution $P_{X}\circ H(\cdot,X_{n+1})$, which generates new “synthetic samples” ${\widetilde{X}}_{n+1}$ which are potentially very similar to this prototype—but coverage may fail for $X_{n+1}$ itself, i.e., marginal coverage may fail to hold, as observed by Guan, (2023).
  • For RLCP, the weights used for computing the threshold $\widehat{q}_{1-\alpha}(X_{n+1},{\widetilde{X}}_{n+1})$ are calculated by centring the kernel $H$ at a random point ${\widetilde{X}}_{n+1}$ that was drawn near $X_{n+1}$, and (as we will see below) will lead to coverage for $X_{n+1}$ itself. We can now think of ${\widetilde{X}}_{n+1}$ as a “synthetic prototype”, and conditionally, the test sample $X_{n+1}$ is likely to lie near to this prototype ${\widetilde{X}}_{n+1}$. More concretely, we will see that the conditional distribution of $X_{n+1}$ given ${\widetilde{X}}_{n+1}$ is equal to the reweighted feature distribution $P_{X}\circ H(\cdot,{\widetilde{X}}_{n+1})$, which (for a strongly localizing kernel $H$) places most of its mass near the “synthetic prototype” ${\widetilde{X}}_{n+1}$.